OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = mx+b y = 3x + b 2 = 3(5) + b solve for b

OpenStudy (anonymous):

Which of the following describes the slopes of perpendicular lines? A. same slope B. opposite slopes C. reciprocal slopes D. opposite reciprocal slopes

OpenStudy (anonymous):

idk how any of these look like

jimthompson5910 (jim_thompson5910):

check out this page http://www.purplemath.com/modules/slope2.htm